OpenStudy (anonymous):

Show that f and g are inverse functions. f(x)=X+3/X-2 g(x)=2x+3/x-1

OpenStudy (anonymous):

f(g(x)) = ? , g(f(x)) = ?

OpenStudy (anonymous):

this algebra is going to stink, but it will be ok

OpenStudy (anonymous):

or you can do algebra as well..

OpenStudy (anonymous):

I tried to solve it but for the answer im not getting x. For inverse function you should always get x

OpenStudy (anonymous):

\[f(g(x))=f(\frac{2x+3}{x-1})\] \[=\frac{\frac{2x+3}{x-1}+3}{\frac{2x+3}{x-1}-2}\] if it is correct, when you simplify this there will be a tone of cancellation and you will get \(x\)

OpenStudy (anonymous):

yes, it is correct http://www.wolframalpha.com/input/?i= \frac{\frac{2x%2B3}{x-1}%2B3}{\frac{2x%2B3}{x-1}-2}

OpenStudy (anonymous):

I have this part done...after this, you multiple 3 by x-1

OpenStudy (anonymous):

3 would have a denominater 1 so it will be 3/1

OpenStudy (anonymous):

multiply top and bottom by \(x-1\) to clear the compound fraction

OpenStudy (anonymous):

you can do this as well : f(x) = (x+3)/(x-2) let f(x) = y y(x-2) = x+3 yx -2y = x+3 yx-x = 2y + 3 x = 2y + 3 / y -1 = g(x)

OpenStudy (anonymous):

you get \[\frac{2x+3+3(x-1)}{2x+3-2(x-1)}\]

OpenStudy (anonymous):

which will give you \(x\) lickety split

OpenStudy (anonymous):

yeh I got 

OpenStudy (anonymous):

sorry its a little crooked

OpenStudy (anonymous):

first step : \[\frac{ \frac{ 2x+3 }{ x-1 } + 3}{ \frac{ 2x+3 }{ x-1 } -2 }\]

OpenStudy (anonymous):

\[\frac{ \frac{ 2x + 3 + 3(x-1) }{ x-1 } }{ \frac{ 2x+3 - 2(x-1) }{ x-1 } }\]

OpenStudy (anonymous):

thanks guyz

OpenStudy (anonymous):

\[\frac{ 2x + 3 + 3x - 3 }{ 2x + 3 -2x + 2 }\] = x